How do you solve the problem x^2=1/9? I tried recriprocals but I am still confused...
Answer by arunpaul(104)

x^2=1/9
0r 9x^2=1
or 9x^2-1=0
or (3x+1)(3x-1) =0
when 3x+1 =0
then 3x =-1
or x = -1/3
when 3x-1 =0
or 3x=1
or x = 1/3
(Ans)